Kansas Company acquired a building valued at $210,000 for property tax purposes in exchange for 12,000 shares of its $5 par common stock. The sKansas Company acquired a building valued at $210,000 for property tax purposes in exchange for 12,000 shares of its $5 par common stock. The stock is widely traded and selling for $15 per share. At what amount should the building be recorded by Kansas Company?

$180,000

$210,000

$120,000

$60,000
Answer #1
Concepts and reason

Assets: It can be defined as the resources owned by the organization which is capable of providing some future benefits. On the basis of duration of time assets are of two types which are 1. Current Assets and 2. Fixed Assets.

Fundamentals

Stock Price: Stock price refers to the market value per share of the shareholder’s equity of the company. The market value of share does not show on the balance sheet of the company.

Preferred Stock: These are the stockholders that are given preference in terms of the dividend over the common stockholders. They receive a fixed amount of dividend as opposed to the common stockholders. They do not carry any right to vote in the general meetings of the company.

Net Income: The resultant amount after reducing all expenses of the company whether direct or indirect for the period from all revenues is termed as net income.

Shareholder’s Equity: It can be referred to as one of the parts of the balance sheet other than the assets and the liabilities of the company. It is also known as owner’s equity of the company. Capital contributed and retained earnings are the parts of the owner’s equity.

Par value: It can be referred to as the face value of the security. Par value of the stock is printed on the face of the stock certificate.

Compute the face value of shares, using the equation as shown below:

Facevalue=Sharesexchanged×Parvalue=12,000shares×$5=$60,000\begin{array}{c}\\{\rm{Face value}} = {\rm{Shares exchanged}} \times {\rm{Par value}}\\\\ = {\rm{12,000 shares}} \times {\rm{\$ 5}}\\\\ = {\rm{\$ 60,000}}\\\end{array}

Hence, the face value of the bond is $60,000.

Compute the value of the building, using the equation as shown below:

Building=Sharesexchanged×Marketvaluepershare=12,000shares×$15=$180,000\begin{array}{c}\\{\rm{Building}} = {\rm{Shares exchanged}} \times {\rm{Market value per share}}\\\\ = {\rm{12,000 shares}} \times {\rm{\$ 15}}\\\\ = {\rm{\$ 180,000}}\\\end{array}

Ans:

The value of the building is $180,000.